# Dense Retrievers Can Fail on Simple Queries: Revealing The Granularity Dilemma of Embeddings

### v1

@@ Abstract
This work focuses on an observed limitation of text encoders: embeddings may not be able to recognize fine-grained entities or events within the semantics, resulting in failed dense retrieval on even simple cases. To examine such behaviors, we first introduce a new evaluation dataset in Chinese, named CapRetrieval, whose passages are image captions, and queries are phrases inquiring entities or events in various forms. Zero-shot evaluation suggests that encoders may fail on these fine-grained matching, regardless of training sources or model sizes. Aiming for enhancement, we proceed to finetune encoders with our proposed data generation strategies, which obtains the best performance on CapRetrieval. Within this process, we further identify an issue of granularity dilemma, a challenge for embeddings to express fine-grained salience while aligning with overall semantics. Our dataset, code and models in this work are publicly released at https://github.com/lxucs/CapRetrieval.
@@
Liyan Xu1     Zhenlin Su2     Mo Yu1     Jiangnan Li1 Fandong Meng1     Jie Zhou1 1Pattern Recognition Center, WeChat AI 2South China University of Technology {liyanlxu,moyumyu}@tencent.com
@@ 1 Introduction
Dense Passage Retrieval (DPR) is a crucial component in search applications, characterized by the dual-encoder paradigm Reimers and Gurevych (2019): queries and candidate passages are independently encoded to text embeddings that capture their semantics, and the system retrieves top candidates based on their embedding similarity. For large language models especially, it serves a pivotal role in Retrieval Augmented Generation (RAG) Lewis et al. (2020). The training of such encoders has become an active direction in representation learning Ramesh Kashyap et al. (2024).
@@
However, despite their growing capabilities encoding complex queries and documents, embedding matching can still fail on simple cases, and yet unable to fully supersede the conventional lexical-based methods such as BM25 Arabzadeh et al. (2021); Ren et al. (2023). Consequently, retrieval systems in practice usually take a hybrid approach for optimized performance Kulkarni et al. (2023); Luo et al. (2023); Chen et al. (2024). In this work, we probe into such limitation of DPR encoders [^1] We focus on single-embedding DPR models in this work. Multi-embedding models, e.g. ColBERT Khattab and Zaharia (2020); Santhanam et al. (2022) and hybrid methods, e.g. M3 Chen et al. (2024) are outside the scope of this study., and show that even for simple passages, the embeddings may often lack a fine-grained view of encoded semantics and the integration of world knowledge, resulting in failed retrieval on entities or events. To remedy this problem, we delve into data generation strategies training for enhanced encoders, and we identify a “granularity dilemma” within this process, posing a challenge to be resolved for general training data composition.
@@
Illustrated by Table 7&8, with strong encoders of both BGE Xiao et al. (2024) and GTE Zhang et al. (2024) series, spanning from 0.1B to 7B models that rank top on the MTEB leaderboard [^2] https://huggingface.co/spaces/mteb/leaderboard for Chinese, all their embeddings fail to rank the more relevant passages against obviously less relevant ones for the query “{CJK*}UTF8gkai炸鸡” (fried chicken) and “{CJK*}UTF8gkai紫色的花” (purple flower) respectively, indicating that such phenomena are universal regardless of training sources and model sizes.
@@
To facilitate the analysis of such behaviors and potential improvements, we first construct a new evaluation dataset in Chinese tailored to a practical image search scenario, where the candidate passages are image captions, and queries are short phrases of entities or events reflected in captions. The dataset overall comprises seemingly simple queries and captions, dubbed CapRetrieval, featuring two unique aspects. First, the retrieval naturally requires fine-grained semantic encoding of passages, inquiring entities or events reflected by the captions in any forms. Second, distinguished from prior datasets related to captions, e.g. Ecom/Video Retrieval Long et al. (2022), which are collected from real-world user clicks but may heavily contain false negatives, our dataset provides explicit annotations for each query-passage pair that enable more reliable and in-depth analysis, constituted by near 1.3 million pair labels.
@@
Upon our zero-shot evaluation on CapRetrieval using open-source encoders of different sizes, all of them are revealed certain flaws performing these fine-grained matching, highlighting the limitations of current embeddings. We then proceed to finetune the encoder with our proposed data generation strategies in Sec. 4, where we leverage LLMs for automatic query generation as training data, yielding full keywords/phrases of passages to enforce the encoding of precise salience within semantics.
@@
Empirical results show that our finetuned 0.1B encoder outperforms all baselines on CapRetrieval, surpassing 7B encoders, validating that the data generation strategies are effective strengthening a more subtle semantic matching. Nevertheless, we also identify a dilemma upon further analysis: with the introduced keywords added into training, the model captures better granular saliency, but may lose grasp on the overall critical semantics in a bigger picture, resulting in degradation on certain scenarios. We describe it as the granularity dilemma to handle both coarse and fine granularity (Sec. 4), showing that it still remains a challenge to obtain general embeddings of a full semantic view.
@@ 2 Dataset: CapRetrieval
A pool of image captions are prepared as candidate passages, derived by three steps. (1) A set of images are collected voluntarily from phone albums of researchers of this work, covering diverse types including photographs, web pictures, and screenshots of various apps or articles. (2) Each image is transcribed into short description (caption) accordingly by prompting GPT-4o. (3) Each caption is manually reviewed and anonymized to ensure privacy compliance. The retrieval only concerns captions, thus the original images are effectively discarded afterwards.
@@
Queries
User queries are short phrases that may be searched to retrieve relevant images, collected by two rounds. (1) The first round of queries are brainstormed by researchers of this work that align with real-world usage scenarios. (2) A combination of queries in the first round is performed, then manually revised to form more complex queries, constituting the second round. Among these queries, the easiest ones can be resolved by lexical matching, while others require fine-grained semantic understanding and world knowledge to resolve.
Labels
Following the convention of typical retrieval datasets in MTEB, the label is a relevance score of $[0,1,2]$ for a pair of query and passage, signaling no/weak/strong relevance. Unlike prior datasets with large-scale candidate pools that prevent annotations for all query-passage pairs, our controlled setting enables full annotation for all pairs by researchers of this work, minimizing false negatives for a reliable and in-depth evaluation.
@@ 3 Zero-Shot Evaluation
To examine the performance of off-the-shelf encoders on CapRetrieval, we evaluate five popular open-source encoder series for Chinese on Huggingface, offering different model sizes:.
@@
• Qwen3777Qwen3-Embedding-{0.6B,8B} Yang et al. (2025): 0.6B / 8B
Our experimental settings comply with the retrieval protocol in MTEB Muennighoff et al. (2023), adopting nDCG@10 as the main evaluation metric. Additionally, since full labels of all pairs are annotated for CapRetrieval, we also provide nDCG@1/5 for more precise evaluation. For each model, we follow the recommended usage by the publishers’ instructions. Performance of BM25 is also provided as a baseline for reference. More implementation details are described in Appx. C.
@@
Table 2: Evaluation results of zero-shot experiments on CapRetrieval, with encoders of different model sizes.
@@ Results
All encoders exhibit flaws matching these fine-grained queries, even the capable 7B model. Though, BM25 still falls short of all encoders already by at least 10%, underscoring that lexical matching is not able to resolve this task.
@@
Table 3 shows the decomposed performance of CapRetrieval by query types (described in B.1). Singleton Entity has the highest nDCG score as 82.05, while Singleton Event and Simple Condition have low scores as 73.2 and 73.8 respectively. The overall trend suggests that the encoder performs worse on more abstract queries, i.e. events or phrases with conditions. Entity-centric queries are relatively easier to resolve (80+ nDCG), though not by a large margin.
@@
The right section of Table 3 presents a comparison between embedding and BM25. BM25 exhibits more polarized performance, where it outperforms embedding on entity-centric queries, but lags behind on more abstract queries. The limitation of BM25 is especially pronounced for Complex Condition, with a 66% gap, highlighting the necessity of using embedding-based retrieval. Overall, it also reveals room for improvement in embeddings as follows.
@@
i) Literal Error: embeddings fail to retrieve passages that contain the full or partial query terms verbatim - passages that BM25 can successfully recall. Though these cases can be remedied by adding lexical search in practice, we advocate that it is important for embeddings to express full information view. Resolving these seemingly simple matches is still challenging for state-of-the-art encoders, which can be regarded as the embedding version of the “needle in a haystack” test for LLMs.
@@
Taxonomy knowledge: certain scenarios require taxonomy involved, e.g. when inquiring a hypernym term such as “household appliances” or “seafood”, or to recognize the matching between “Cantonese-style roasted meats” and “char siu”.
@@
Ignoring subjects or conditions: the passage only addresses partial semantics but fails to accommodate full conditions. For instance, “purple flower” retrieves captions about “purple butterflies”; “shopping cart screenshot” retrieves a screenshot but of a ride-hailing app. As this kind of errors are quite common for queries with conditions, it suggests that current embeddings may not actually encode “concepts”, but in a way more towards superficial matching.
@@ 4 Encoder Training
The results of zero-shot evaluation on CapRetrieval calls for more expressive embeddings that capture fine-grained semantics and world knowledge integration. We proceed further examination by finetuning encoders with sets of small-scale experiments.
@@ 4.1 Training Data Generation
Training pairs in large-scale retrieval resources such as mMARCO Bonifacio et al. (2022) and DuReader Qiu et al. (2022) mainly comprise user search queries and clicks. Consequently, for a passage, the queries associated with a given passage in the training set are often coarse-grained, such that they do not address the full semantic content. Motivated towards fine-grained semantic matching, we propose automatic query generation for enhanced training, with distinct granularity as follows.
@@
  • In-domain (ID): we collect more image captions as the training passage pool. To mitigate memorization, we filter out all captions of ROUGE-L $>0.6$ w.r.t. any test captions in CapRetrieval.
@@
Passage: {CJK*}UTF8gkai图片显示了上海市电力公司的月度账单,包括2021年5月至9月的用电费用和支付状态。 (The image shows the monthly bill from the Shanghai Electric Power Company, including electricity charges and payment status from May to September 2021.)
@@
Bill Electricity fee Utility bill Payment status Monthly electricity charges Utility payment screenshot 2021 electricity fee Shanghai Electric Power Company bill inquiry Electricity usage details from May to September 2021 Monthly electricity bill details Electricity payment status records Historical payment information from Shanghai Electric Power Company Shanghai Electric Power Company electricity bill from May to September 2021
@@ 4.2 Granularity Dilemma
Upon further analysis, we identified that though training with keywords could facilitate fine-grained matching, it may overlook the overall saliency. For example, the query in VideoRetrieval “{CJK*}UTF8gkai荒野独居第2季中文版” (Alone Season 2 Chinese Edition) should retrieve the TV show Alone with the specific requirement; whereas the encoder trained with KW may overemphasize terms on Season 2 or Chinese, but fails to correctly prioritize the supposedly most critical concept Alone, showing a misalignment of semantic importance. We refer to this observation as the granularity dilemma.
@@ 5 Conclusion
In this work, we focus on the granularity issue of embeddings, stemming from the observation that encoders can fail on entity or event matching of even simple cases. A new evaluation set in Chinese is introduced to probe the limitation of embeddings, and zero-shot experiments reveal the room for improvement on these fine-grained matching. To remedy this problem, we investigate automatic data generation strategies utilizing LLMs, aiming for strengthening the full semantic view of embeddings through additional training. As our trained encoder of 0.1B surpasses the best baseline of 7B, we further identify the granularity dilemma within this process, which calls for future efforts on the granularity issue towards expressive embeddings.
@@
No external annotators were recruited or employed during the dataset creation process; all annotation and verification were conducted internally by in-house researchers of this work (two males with the native language as Chinese). The data preparation and annotation process is approved and audited by the in-house research department. All passages are derived from image transcriptions generated from GPT-4o. As such, there are no visible risks, copyright or legal concerns in using this dataset.
ChatGPT from OpenAI is used to assist the writing of this paper.
@@ Appendix A Related Work
Embedding-based text representation has been a fundamental direction in representation learning since the advent of BERT-style encoding Devlin et al. (2019), and plays a critical role in modern NLP applications including various retrieval tasks. Early works such as S-BERT Reimers and Gurevych (2019), SimCSE Gao et al. (2021) and Contriever Izacard et al. (2022) establish the effective training paradigm of contrastive learning on unsupervised or weakly supervised text pairs. Current state-of-the-art encoders usually adopt a multi-stage training that consists of both unsupervised and supervised finetuning stages Neelakantan et al. (2022); Wang et al. (2024a); Li et al. (2023b); Xiao et al. (2024). Among the popular supervised training resources, most of the datasets are collected through real-world user behaviors, such as MSMARCO Nguyen et al. (2017) and DuReader He et al. (2018); Qiu et al. (2022). Recently, synthetic data generation by LLMs is also reported positive gains in encoder training Li et al. (2024); Wang et al. (2024b); Yang et al. (2025).
Beyond the conventional single-embedding encoders, other paradigms have been proposed for retrieval, such as ColBERT with token-level embeddings Khattab and Zaharia (2020); Santhanam et al. (2022), hybrid encoders with lexical features Kulkarni et al. (2023); Luo et al. (2023) and sparse features Chen et al. (2024).
The analysis on embedding dynamics and characteristics is also an active research direction, such as the embedding geometry Reif et al. (2019); Garí Soler and Apidianaki (2021) and embedding inversion Li et al. (2023a); Morris et al. (2023).

### v2

@@ Abstract
This work stems from an observed limitation of text encoders: embeddings may not be able to recognize fine-grained entities or events within encoded semantics, resulting in failed retrieval even in simple cases. To examine such behaviors, we first introduce a new evaluation dataset, CapRetrieval, in which passages are image captions and queries are phrases targeting entity or event concepts in diverse forms. Zero-shot evaluation suggests that encoders often struggle with these fine-grained matching, regardless of training sources or model size. Aiming for enhancement, we proceed to finetune encoders with our proposed data generation strategies, enabling a small 0.1B encoder to outperform the state-of-the-art 7B model. Within this process, we further uncover the granularity dilemma, a challenge for embeddings to capture fine-grained salience while aligning with overall semantics. Our dataset, code and models in this work are publicly released at https://github.com/lxucs/CapRetrieval.
@@
Liyan Xu1    Zhenlin Su2    Mo Yu1    Jiangnan Li1 Fandong Meng1     Jie Zhou1 1Pattern Recognition Center, WeChat AI 2South China University of Technology liyanlxu@tencent.com
@@ 1 Introduction
Dense Passage Retrieval (DPR) is a crucial component in searching, characterized by the classic dual-encoder paradigm Reimers and Gurevych (2019): queries and candidate passages are independently encoded to embeddings that capture their semantics, then top candidates are retrieved based on their embedding similarity. For large language models (LLMs) especially, DPR serves a pivotal role in Retrieval Augmented Generation (RAG) Lewis et al. (2020). The training of such encoders has thereby become an active direction in representation learning Ramesh Kashyap et al. (2024).
@@
However, despite their growing capabilities in encoding complex queries and documents, embedding matching can still fail on simple cases, and yet unable to fully supersede the conventional lexical-based methods such as BM25 Arabzadeh et al. (2021); Ren et al. (2023). Consequently, retrieval systems in practice usually take a hybrid approach for optimized performance Chen et al. (2022); Kulkarni et al. (2023); Luo et al. (2023); Chen et al. (2024). In this work, we probe into such limitation of DPR encoders [^1] We focus on embedding-based DPR in this work, and leave reranking or other hybrid paradigms as a separate scope., and highlight that even for simple passages, state-of-the-art encoders still often lack a fine-grained view of concepts and the integration of world knowledge, resulting in failed retrieval on entities or events. To remedy this issue, we delve into data generation strategies towards enhanced encoder training, and we identify a granularity dilemma within this process, posing a challenge for general training data composition.
@@
Illustrated by Table 8&9, with strong encoders of both BGE Xiao et al. (2024) and GTE Zhang et al. (2024) series, spanning from 0.1B to 7B models that rank top on the MTEB leaderboard [^2] https://huggingface.co/spaces/mteb/leaderboard for Chinese, all their embeddings fail to rank the more relevant passages against obviously less relevant ones, for the simple query fried chicken, purple flower, and watermelon respectively, indicating that such phenomena occur universally, regardless of training sources, model sizes or languages.
@@
To facilitate the analysis of such behaviors and potential improvements, we first construct a new evaluation dataset in both Chinese and English, tailored to a practical image search scenario, where the candidate passages are image captions, and queries are short phrases of entities or events reflected in captions. The dataset overall comprises seemingly straightforward queries and captions, dubbed CapRetrieval, and features two unique aspects. First, it naturally requires fine-grained semantic encoding of passages, inquiring concepts of entities or events reflected in the captions. Second, distinguished from prior caption-related datasets, e.g. Ecom/Video Retrieval Long et al. (2022), whose labels are collected from real-world user clicks and thus may contain many false negatives, CapRetrieval provides explicit annotations for each query-passage pair, enabling more reliable and in-depth analysis by near 1.3 million pair labels.
@@
Upon our zero-shot evaluation on CapRetrieval using open-source encoders of various sizes, all of them are revealed certain flaws performing these fine-grained matching, underscoring the current limitations of embeddings. We then proceed to finetune the encoder with our proposed data generation strategies in Sec. 4, where we leverage LLMs for automatic keyword generation as training data to enforce the semantic encoding of precise salience.
@@
Empirical results show that our finetuned 0.1B encoder outperforms all baselines on CapRetrieval, surpassing 7B encoders, validating that the data generation strategies are effective strengthening a more subtle semantic matching. Nevertheless, we also identify a dilemma upon further analysis: with the introduced keywords added into training, the model improves on granular saliency and concepts, but may lose grasp on the overall critical semantics in a bigger picture, resulting in degradation on certain scenarios. We describe it as the granularity dilemma to handle both coarse and fine granularity (Sec. 4), showing that it still remains a challenge to obtain general embeddings of a full semantic view.
@@ 2 Dataset: CapRetrieval
A pool of image captions are prepared as candidate passages, derived from three steps. (1) A set of images is collected from personal phone albums for diverse coverage of different types, including daily photographs, web pictures, screenshots of various apps or articles. (2) Each image is transcribed into a short caption by prompting GPT-4o. (3) Each caption is manually reviewed and anonymized to ensure privacy compliance.
Queries
User queries are short phrases to search relevant images, collected by two rounds. The first round of queries are brainstormed by researchers of this work, being usual entities or events that align with real-world usage scenarios. For the second round, a combination of queries in the first round is performed, then manually revised to form more complex queries. Among these queries, the easiest ones can be resolved by lexical matching, while others require fine-grained concept understanding and world knowledge to resolve.
Labels
Following the convention of typical retrieval datasets in MTEB, the label is a relevance score of $[0,1,2]$ for a pair of query and passage, signaling no/weak/strong relevance. Unlike prior web-collected datasets with large-scale passage pools, our controlled setting enables exhaustive annotations for each query-passage pair, which minimizes false negatives for a more reliable evaluation.
Note that the annotation and the entire retrieval process only concern textual captions, without involving vision features. The accuracy of image captioning thus does not affect the evaluation.
@@
CapRetrieval was originally constructed in Chinese. We further employ GPT-4o to translate all queries and passages into English to provide the corresponding English version, of which the translation quality is verified by human experts. The remaining sections in this paper reports on the original CapRetrieval without further specification.
@@ 3 Zero-Shot Evaluation
To examine the performance of off-the-shelf encoders on CapRetrieval, we evaluate five popular open-source encoder series for Chinese on Huggingface, offering various model sizes:
@@
• Qwen3777Qwen3-Embedding-{0.6B,8B} Zhang et al. (2025): 0.6B / 8B
Our experimental settings comply with the retrieval protocol in MTEB Muennighoff et al. (2023), adopting nDCG@10 as the main evaluation metric. Additionally, since full labels of all pairs are annotated for CapRetrieval, we also provide nDCG@1/5 for more precise evaluation. For each model, we follow the recommended usage by its publishers’ instructions. Performance of BM25 is also provided as a baseline for reference. More implementation details are described in Appx. C.
@@
Human 100.00 98.57 97.83
Table 2: Evaluation results of zero-shot experiments on CapRetrieval, with encoders of different model sizes. Human performance is evaluated on a 10% subset. Evaluation on the according English dataset (CapRetrievalEn) is separately presented in Table 6.
@@ Results
All encoders exhibit flaws matching these fine-grained queries, even the capable 7B/8B LLM models. Though, they all outperform BM25 by at least 10%, underscoring that lexical matching alone is not able to resolve this task.
@@
Table 3 shows the decomposed performance on CapRetrieval by query types (described in B.1). Singleton Entity has the highest nDCG score as 82.05, while Singleton Event and Simple Condition have low scores as 73.2 and 73.8 respectively. The overall trend suggests that the encoder performs worse on more abstract queries, i.e. events or phrases with conditions. Entity-centric queries are relatively easier to resolve (80+ nDCG), though not by a large margin.
@@
The right section of Table 3 presents a comparison between embedding and BM25. BM25 exhibits more polarized performance, where it outperforms embedding on entity-centric queries, but lags behind on more abstract queries. The limitation of BM25 is especially pronounced for Complex Condition, with a 66% gap, highlighting the necessity of embedding-based retrieval. Overall, it also reveals room for improvement in embeddings as follows.
@@
i) Literal Error: embeddings fail to retrieve passages that contain the full or partial query terms verbatim - passages that BM25 can successfully recall. Though these cases can be remedied by adding lexical search in practice, we advocate that embeddings should encode concepts with full information view. Resolving these seemingly simple matches is still challenging for state-of-the-art encoders, which can be regarded as the embedding version of the “needle in a haystack” test for LLMs.
@@
Taxonomy knowledge: certain scenarios require taxonomy knowledge involved, e.g. when inquiring a hypernym term such as “household appliances” or “seafood”, or to recognize the matching between “Cantonese-style roasted meats” and “char siu”.
@@
Ignoring subjects or conditions: the passage only addresses partial semantics but fails to accommodate full concepts, e.g. “purple flower” retrieves captions about purple butterflies; “shopping cart screenshot” retrieves a screenshot but of a ride-hailing app. As this kind of errors are quite common for queries with conditions, it suggests that embeddings may not actually encode concepts, but in a way towards superficial matching.
@@ 4 Encoder Training
The results of zero-shot evaluation on CapRetrieval calls for more expressive embeddings that capture fine-grained concepts and world knowledge integration. We proceed with further examination by finetuning encoders with training data strategies.
@@ 4.1 Training Data Generation
Training pairs in existing large-scale resources such as mMARCO Bonifacio et al. (2022) and DuReader Qiu et al. (2022) mainly comprise user search queries and clicks. Consequently, for a passage, the queries associated with a given passage in the training set are often coarse-grained, such that they do not address the full semantic content. Motivated towards fine-grained semantic matching, we propose automatic query generation for enhanced training, with distinct granularity as follows.
@@
  • In-domain (ID): we collect more image captions as the training passage pool. To mitigate memorization, we filter out all captions of ROUGE-L >0.6>0.6 w.r.t. any test captions in CapRetrieval.
@@
Passage: 图片显示了上海市电力公司的月度账单,包括2021年5月至9月的用电费用和支付状态。
(The image shows the monthly bill from the Shanghai Electric Power Company,
including electricity charges and payment status from May to September 2021.)
@@
Bill
Electricity fee
Utility bill
Payment status
Monthly electricity charges
Utility payment screenshot
2021 electricity fee
Shanghai Electric Power Company bill inquiry
Electricity usage details from May to September 2021
Monthly electricity bill details
Electricity payment status records
Historical payment information from Shanghai Electric Power Company
Shanghai Electric Power Company electricity bill from May to September 2021
@@ 4.2 The Granularity Dilemma
Upon further analysis, we identified that though training with keywords could facilitate fine-grained matching, it may overlook the overall saliency. For example, the query in VideoRetrieval “荒野独居第2季中文版” (Alone Season 2 Chinese Edition) should retrieve the TV show Alone with the specified requirement; whereas the encoder trained with KW may overemphasize terms on Season 2 or Chinese, but fails to correctly prioritize the supposedly most critical concept Alone, showing a misalignment of semantic importance. We refer to this observation as the granularity dilemma.
@@ 5 Conclusion
We focus on the embedding granularity that stems from the observation, where text encoders can fail to recognize entities or events of even simple cases. A new evaluation set is introduced to probe the limitation of embeddings, and zero-shot experiments reveal the room for improvement on these fine-grained matching. We further investigate data generation strategies for encoder training, and identify the granularity dilemma that calls for future efforts towards more expressive embeddings.
@@
No external annotators were recruited or employed during the dataset creation process; all annotation and verification were conducted internally by in-house researchers of this work. The data preparation and annotation process is approved and audited by the in-house research department. All passages are derived from image transcriptions generated from GPT-4o. As such, there are no visible risks, copyright or legal concerns in using this dataset.
@@ Appendix A Related Work
Dense retrieval serves a critical role and has received growing attention in recent developments of Retrieval-Augmented Generation (RAG) Lewis et al. (2020); Sarthi et al. (2024); Xu et al. (2024); Edge et al. (2025); Qian et al. (2025); Wang et al. (2025). As a fundamental direction in representation learning, early works such as S-BERT Reimers and Gurevych (2019), SimCSE Gao et al. (2021) and Contriever Izacard et al. (2022) establish the effective training paradigm for embedding-based text representation, using contrastive learning on unsupervised or weakly supervised text pairs. Current state-of-the-art encoders usually adopt a multi-stage training that consists of both unsupervised and supervised finetuning stages Neelakantan et al. (2022); Wang et al. (2024a); Li et al. (2023); Xiao et al. (2024). Among the popular supervised training resources, most of the datasets are collected through real-world user behaviors, such as MSMARCO Nguyen et al. (2017) and DuReader He et al. (2018); Qiu et al. (2022). Recently, synthetic data generation by LLMs is also reported positive gains in encoder training Li et al. (2024); Wang et al. (2024b); Yang et al. (2025).
Beyond the conventional single-embedding encoders, other paradigms have been proposed for retrieval, such as ColBERT with token-level embeddings Khattab and Zaharia (2020); Santhanam et al. (2022), hybrid encoders with lexical features Chen et al. (2022); Kulkarni et al. (2023); Luo et al. (2023) and sparse features Chen et al. (2024). Recent efforts have also explored connecting global dependencies within embeddings Wu et al. (2025).